# How Muon’s Spectral Design Benefits Generalization: A Study on Imbalanced Data

### v1

@@ 3.3 Effect of Depth
We begin with the case $L=2$, which is particularly instructive due to its connection to the Unconstrained Features Model (UFM). The UFM is widely used to study the geometry of learned representations in deep and nonlinear networks under finite-sample regimes (Yang et al.,, 2018; Mixon et al.,, 2020). It serves as a proxy for architectures with sufficient capacity to freely optimize both the feature representation ($\bm{W}^{0}\bm{x}$) and the classification head ($\bm{W}^{1}$). The bilinear model recovers this framework when the input features $\bm{x}$} are orthogonal, making it a natural setting for analyzing the interplay between feature learning and classifier dynamics.
@@
Suppose Condition 1 holds. Let the weights be initialized as $\bm{W}^{0}_{0}=e^{-\delta}\bm{Q}[{\bm{I}}_{d_{1}}\,\,\mathbf{0}_{d-d_{1}}]{\bm{V}}^{\top}$ and $\bm{W}^{1}_{0}=e^{-\delta}{\bm{U}}[{\bm{I}}_{k}\,\,\mathbf{0}_{d_{1}-k}]\bm{Q}^{\top}$, where $\bm{Q}\in\mathbb{R}^{d_{1}\times d_{1}}$ is an orthonormal matrix, $k<d_{1}<d$, and $\delta>0$ is a constant. Then, for SpecGD, at each iteration, $\bm{W}^{0}_{t}:=\bm{Q}\overline{\bm{W}}^{0}_{t}{\bm{V}}^{\top}$, $\bm{W}_{t}^{1}={\bm{U}}\overline{\bm{W}}^{1}_{t}\bm{Q}^{\top}$, where $\overline{\bm{W}}^{0}_{t},\overline{\bm{W}}^{1}_{t}$ are zero except their main diagonal along which, in the limit $\delta\rightarrow\infty$, they evolve as follows for $c\in[k]$:
@@
See Sec.˜B.5 for the full version and the proof. We find that, in contrast to the linear model where SpecGD learns the $c^{\text{th}}$ component in time $t_{c}=\tfrac{1}{\eta}\tfrac{s^{\texttt{yx}}_{c}}{s^{\texttt{xx}}_{c}}$, for the bilinear model, it saturates in time $t_{c}\approx\tfrac{1}{\eta}\sqrt{\tfrac{s^{\texttt{yx}}_{c}}{s^{\texttt{xx}}_{c}}}$ (for large $\delta$). This is quite different from the effect of increasing model depth on the dynamics of GD. As discussed in Saxe et al., (2013); Gidel et al., (2019), for GD, the time to reach $\epsilon$-close to the optimal (for the $c^{\text{th}}$ component) is $t_{c}=\tfrac{\ln(\epsilon)}{\ln(1-\eta s^{\texttt{xx}}_{c})}\approx\tfrac{\ln(1/\epsilon)}{\eta s^{\texttt{xx}}_{c}}$ (for small $\eta$) for a linear model, compared to $t_{c}\approx\tfrac{\ln\left(1/\epsilon-1\right)+\ln\left(s^{\texttt{xx}}_{c}e^{\delta}/s^{\texttt{yx}}_{c}-1\right)}{2\eta s^{\texttt{yx}}_{c}}\approx\tfrac{\ln(1/\epsilon)+\delta}{2\eta s^{\texttt{yx}}_{c}}$ (for small $\epsilon$ and large $\delta$) for a bilinear model. To make the comparison more concrete, recall from Lemma˜3.1 that for our class-imbalanced setting with data model (DM), when $\mu=1$, $s^{\texttt{xx}}_{c}=p_{c}+\texttt{\text{SNR}}^{-1}$ and $s^{\texttt{yx}}_{c}=p_{c}$. The GD dynamics for linear vs. bilinear model depend on the SNR: when SNR is high, we observe the stage-wise learning dynamics in both cases, i.e., spectral components are learned in descending order of class priors. However, when SNR is low, these stage-wise dynamics may not be apparent for the linear model as $t_{c}\propto\tfrac{1}{p_{c}+\texttt{\text{SNR}}^{-1}}\approx\texttt{\text{SNR}}$. In contrast, for SpecGD, $t_{c}\propto\left(\tfrac{p_{c}}{p_{c}+\texttt{\text{SNR}}^{-1}}\right)^{1/L}$, where $L\in\{1,2\}$. This means the dynamics are qualitatively similar for both models; however, the gap between saturation times on different components is smaller for the deeper bilinear model.
@@
We next consider deeper models with L>2L\!>\!2. The dynamics of $\bm{W}_{t}=\prod_{i=0}^{L-1}\bm{W}_{t}^{L-1-i}$ for SpecGD can be characterized in a similar way as Proposition˜2. Specifically, $\{\overline{\bm{W}}^{l}_{t}\}_{l=0}^{L-1}$ defined analogously remain diagonal with entries $c\in[k]$, in the limit $\delta\!\rightarrow\!\infty$, evolving as:
@@
We find that as $L$ increases, the gap between the saturation of the different components becomes smaller. We empirically validate this result for SpecGD in Fig.˜6 by comparing the test accuracy dynamics of $2$-layer and $4$-layer MLPs on the Colored-MNIST dataset. A similar trend is observed for NGD, where depth speeds up learning of the minority component. This is an interesting empirical parallel, given that precise theoretical dynamics for GD have primarily been characterized for depth $L\leq 2$ (Gidel et al.,, 2019; Saxe et al.,, 2013).
@@ 4 Experimental Results
In this section, we present experimental results on datasets with group and class imbalance, comparing the performance of Muon to SGD. Our analysis is guided by the spectral perspective developed in Sec.˜3.2. For class imbalance, recall that spectral components correspond to class priors. We extend this analogy to group imbalance, where inputs contain both core and spurious features, and groups are defined by different core-spurious feature combinations. As seen in our motivating Colored-MNIST example (Fig.˜1), spurious features (e.g., color) act as the dominant spectral component, while core features (e.g., digit shape) are less dominant (see also (Ng et al.,, 2024), which shows demonstrates this for some other datasets with spurious correlations). Our experiments show that Muon promotes a more balanced learning of different spectral components, thereby improving generalization relative to SGD. Finally, while our theory focused on the dynamics of SpecGD versus (N)GD, our empirical study also includes Adam, offering a broader comparison with a widely used adaptive optimizer.
@@
Dominoes Dataset. We consider the MNIST-CIFAR dataset from the Dominoes benchmark (Shah et al.,, 2020; Pagliardini et al.,, 2022), where each image is generated by stacking an MNIST digit from class $\{0,1\}$ on the top with a CIFAR-10 image from class $\{\text{automobile, truck}\}$ on the bottom. The MNIST part is spuriously correlated with the label $95\%$ of the time, while the CIFAR part is $100\%$ predictive of the label. Example images are shown in Fig.˜7 (left).
@@
Subgroup Robustness Benchmarks. Next, we consider two benchmark subgroup robustness datasets: MultiNLI (Williams et al.,, 2018) and CelebA (Liu et al.,, 2015). MultiNLI consists of sentence pairs, where the task is to classify the relationship of the second sentence to the first as entailment, neutral, or contradiction. The spurious feature is the presence of negation words, which often indicate contradiction. CelebA contains images of celebrity faces, with the task of predicting whether the hair color is blonde/not blonde, and the celebrity’s gender (male/female) serves as the spurious attribute.
@@
We fine-tune a pretrained BERT bert-base-uncased model (Devlin et al.,, 2019) on the MultiNLI dataset, and an ImageNet-pretrained ResNet-50 model on CelebA, using different optimizers and evaluate the average (group-balanced) test accuracy and the worst-group accuracy over the number of finetuning epochs. To ensure our comparisons are robust, we conduct extensive hyperparameter sweeps for all optimizers (see Sec.˜D.4 for details). We select the best hyperparameter configuration for any given optimizer and number of fine-tuning epochs, $T$, based on the highest average worst-group accuracy achieved on the validation set, averaged across multiple initialization seeds, at any point up to $T$ epochs. We then report the performance for the optimal configuration for different FT epochs. From Fig.˜8, we see that Muon outperforms SGD on both datasets. On MultiNLI, Adam attains better performance than Muon/Shampoo whereas on CelebA, its performance is comparable to Muon. These experiments show that Muon promotes more balanced learning of spectral components than SGD, thereby improving generalization.
@@
We now extend our investigation of class imbalance to language modeling. While our previous experiments and theory focused on one-hot classification, we now shift our comparison of optimizers to a next-token prediction (NTP) task. This task presents a natural imbalance over the distribution of tokens, as word frequencies follow a long-tail distribution described by Zipf’s law (Piantadosi,, 2014).
@@
TinyStories Dataset. To examine how different optimizers compare on rare token learning, we train small Transformer models on the TinyStories (Eldan and Li,, 2023) corpus and compare the performance of Muon, SGD, and Adam. We group tokens into "frequent" and "rare" categories based on their corpus statistics and evaluate the token-level accuracy for each group separately on a held-out test set. Complete training details are available in App. D.5. The performance on rare and frequent tokens is reported in Fig.˜9. Consistent with our findings in the one-hot class-imbalanced classification, we observe that Muon generalizes faster on both the majority (frequent) and minority (rare) components compared to SGD, and in this case, even compared to Adam.
Attribute-Organism Classification. To extend our findings from statistical imbalance to a setting with explicit conceptual structure, we conduct a toy experiment on hierarchical concept learning (Zhao and Thrampoulidis,, 2025). In this task, a model must classify organisms (e.g., "dog") based on their attributes, which are shared at different levels of a taxonomic tree (e.g., "needs oxygen" for all animals). We observe distinct learning dynamics: GD learns the coarse, high-level categories (plant vs. animal) first, while Muon exhibits a more balanced learning progression across both coarse and fine-grained concepts (see Sec.˜D.6 for details)
@@ 5 Discussion and Outlook
A crucial test of this insight is how its practical variants, Muon and Shampoo, perform against strong, widely-used baselines like Adam. In our experiments in Sec.˜4, we find that Adam’s generalization performance is only slightly worse—or sometimes even better—than that of Muon and Shampoo. This is consistent with recent large-scale benchmarks for LLM pretraining, which show that well-tuned Adam variants remain highly competitive (Wen et al.,, 2025; Semenov et al.,, 2025). Previous work has used different perspectives to explain Adam’s superiority over (S)GD: i) Kunstner et al., (2024) show how heavy-tailed class-imbalanced settings lead to gains for Adam by focusing on their optimization dynamics; ii) Vasudeva et al., 2025b show that Adam promotes richer feature learning than SGD in neural networks, which can lead to better generalization in group-imbalanced settings. Together, these works underscore Adam’s advantages in imbalanced regimes, albeit through different mechanisms. Our motivation for using this testbed stems from recognizing that spectral components of imbalanced data directly correspond to majority and minority classes or groups, and our analysis reveals that SpecGD achieves its gains through a fundamentally different and interpretable mechanism—balanced learning of principal components.
A natural question is how the implicit balancing effect of spectral methods compares to explicitly designed techniques like loss re-weighting (Byrd and Lipton,, 2019; Cao et al.,, 2019; Sagawa et al.,, 2019; Menon et al.,, 2020; Liu et al.,, 2021). In Sec.˜D.7, we show that while weighted cross-entropy (wCE) effectively encourages balanced learning of spectral components on NMD, Muon provides a similar, though weaker, effect "for free" without requiring explicit group priors. This suggests spectral methods could be useful when such information is unavailable for re-weighting.
Our work provides a framework to understand the generalization of spectrum-aware optimizers within a controlled setting using population statistics and squared loss. A natural next step is to build upon this foundation by extending the analysis to more complex finite-sample and cross-entropy regimes. We outline these exciting future directions and their associated challenges in App.˜F.
@@ Appendix A Optimizers
Using this, we have $\mathbf{\Delta}_{t}=\bm{L}_{t}^{-1/4}\bm{M}_{t}\bm{R}_{t}^{-1/4}$. It is easy to see that Shampoo reduces to SpecGD when we set $\beta_{2}=0,\beta_{1}=0$.
@@
$\mathbf{\Delta}_{t}=(\hat{\bm{Z}}_{t}+\epsilon\bm{1}\bm{1}^{\top})^{\circ-1/2}\odot\hat{\bm{M}}_{t}$, where $(\cdot)^{\circ}$ denotes the Hadamard power, $\epsilon>0$ is the numerical precision parameter, and $\mathbf{1}$ denotes the all-ones vector. It is easy to see that Adam reduces to SignGD when we set $\beta_{1}=\beta_{2}=\epsilon=0$.
@@ B.2 Proof of Proposition˜1
The continuous–time dynamics are
@@
The dynamics take the form

### v2

@@ 3.3 Effect of Depth
We begin with the case $L=2$, which is particularly instructive due to its connection to the Unconstrained Features Model (UFM). The UFM is widely used to study the geometry of learned representations in deep and nonlinear networks under finite-sample regimes (yang2018breakingsoftmaxbottleneckhighrank; mixon2020neural). It serves as a proxy for architectures with sufficient capacity to freely optimize both the feature representation ($\bm{W}^{0}\bm{x}$) and the classification head ($\bm{W}^{1}$). The bilinear model recovers this framework when the input features $\bm{x}$} are orthogonal, making it a natural setting for analyzing the interplay between feature learning and classifier dynamics.
@@
Suppose Condition 1 holds. Let the weights be initialized as $\bm{W}^{0}_{0}=e^{-\delta}\bm{Q}[{\bm{I}}_{d_{1}}\,\,\mathbf{0}_{d-d_{1}}]{\bm{V}}^{\top}$ and $\bm{W}^{1}_{0}=e^{-\delta}{\bm{U}}[{\bm{I}}_{k}\,\,\mathbf{0}_{d_{1}-k}]\bm{Q}^{\top}$, where $\bm{Q}\in\mathbb{R}^{d_{1}\times d_{1}}$ is an orthonormal matrix, $k<d_{1}<d$, and $\delta>0$ is a constant. Then, for SpecGD with, at each iteration, $\bm{W}^{0}_{t}:=\bm{Q}\overline{\bm{W}}^{0}_{t}{\bm{V}}^{\top}$, $\bm{W}_{t}^{1}={\bm{U}}\overline{\bm{W}}^{1}_{t}\bm{Q}^{\top}$, where $\overline{\bm{W}}^{0}_{t},\overline{\bm{W}}^{1}_{t}$ are zero except their main diagonal along which, in the limit $\delta\rightarrow\infty$, they evolve as follows for $c\in[k]$:
@@
For clarity, we consider $\delta\to\infty$ in the above proposition to state the SpecGD dynamics. See Sec.˜B.5 for the complete version (for any $\delta>0$) and the proof.
We find that, in contrast to the linear model where SpecGD learns the $c^{\text{th}}$ component in time $t_{c}=\tfrac{1}{\eta}\tfrac{s^{\texttt{yx}}_{c}}{s^{\texttt{xx}}_{c}}$, for the bilinear model, it saturates in time $t_{c}\approx\tfrac{1}{\eta}\sqrt{\tfrac{s^{\texttt{yx}}_{c}}{s^{\texttt{xx}}_{c}}}$ (for large $\delta$). This is quite different from the effect of increasing model depth on the dynamics of GD. As discussed in saxe2013exact; gidel2019implicit, for GD, the time to reach $\epsilon$-close to the optimal (for the $c^{\text{th}}$ component) is $t_{c}=\tfrac{\ln(\epsilon)}{\ln(1-\eta s^{\texttt{xx}}_{c})}\approx\tfrac{\ln(1/\epsilon)}{\eta s^{\texttt{xx}}_{c}}$ (for small $\eta$) for a linear model, compared to $t_{c}\approx\tfrac{\ln\left(1/\epsilon-1\right)+\ln\left(s^{\texttt{xx}}_{c}e^{\delta}/s^{\texttt{yx}}_{c}-1\right)}{2\eta s^{\texttt{yx}}_{c}}\approx\tfrac{\ln(1/\epsilon)+\delta}{2\eta s^{\texttt{yx}}_{c}}$ (for small $\epsilon$ and large $\delta$) for a bilinear model. To make the comparison more concrete, recall from Lemma˜3.1 that for our class-imbalanced setting with data model (DM), when $\mu=1$, $s^{\texttt{xx}}_{c}=p_{c}+\texttt{\text{SNR}}^{-1}$ and $s^{\texttt{yx}}_{c}=p_{c}$. The GD dynamics for linear vs. bilinear model depend on the SNR: when SNR is high, we observe the stage-wise learning dynamics in both cases, i.e., spectral components are learned in descending order of class priors. However, when SNR is low, these stage-wise dynamics may not be apparent for the linear model as $t_{c}\propto\tfrac{1}{p_{c}+\texttt{\text{SNR}}^{-1}}\approx\texttt{\text{SNR}}$. In contrast, for SpecGD, $t_{c}\propto\left(\tfrac{p_{c}}{p_{c}+\texttt{\text{SNR}}^{-1}}\right)^{1/L}$, where $L\in\{1,2\}$. This means the dynamics are qualitatively similar for both models; however, the gap between saturation times on different components is smaller for the deeper bilinear model.
@@
We next consider deeper models with $L\!\geq\!2$. The dynamics of $\bm{W}_{t}=\prod_{i=0}^{L-1}\bm{W}_{t}^{L-1-i}$ for SpecGD can be characterized in a similar way as Proposition˜2. Specifically, $\{\overline{\bm{W}}^{l}_{t}\}_{l=0}^{L-1}$ defined analogously remain diagonal with entries $c\in[k]$, in the limit $\delta\!\rightarrow\!\infty$, evolving as:
@@
We find that as $L$ increases, the gap between the saturation of the different components becomes smaller. We empirically validate this result for SpecGD in Fig.˜6 by comparing the test accuracy dynamics of $2$-layer and $4$-layer MLPs on the Colored-MNIST dataset. A similar trend is observed for NGD, where depth speeds up learning of the minority component. This is an interesting empirical parallel, given that precise theoretical dynamics for GD have primarily been characterized for depth $L=2$ (gidel2019implicit; saxe2013exact).
@@ 4 Experimental Results
In this section, we present experimental results on datasets with group and class imbalance, comparing the performance of Muon to SGD. Our analysis is guided by the spectral perspective developed in Sec.˜3.2. For class imbalance, recall that spectral components correspond to class priors. We extend this analogy to group imbalance, where inputs contain both core and spurious features, and groups are defined by different core-spurious feature combinations. As seen in our motivating Colored-MNIST example (Fig.˜1), spurious features (e.g., color) act as the dominant spectral component, while core features (e.g., digit shape) are less dominant (see also (ng2024understanding), which demonstrates this for some other datasets with spurious correlations). Our experiments show that Muon promotes a more balanced learning of different spectral components, thereby improving generalization relative to SGD. Finally, while our theory focused on the dynamics of SpecGD versus (N)GD, our empirical study also includes Adam, offering a broader comparison with a widely used adaptive optimizer.
@@
Dominoes Dataset. We consider the MNIST-CIFAR dataset from the Dominoes benchmark (shah2020pitfalls; pagliardini2022agreedisagreediversitydisagreement), where each image is generated by stacking an MNIST digit from class $\{0,1\}$ on the top with a CIFAR-10 image from class $\{\text{automobile, truck}\}$ on the bottom. The MNIST part is spuriously correlated with the label $95\%$ of the time, while the CIFAR part is $100\%$ predictive of the label. Example images are shown in Fig.˜7 (left).
@@
Subgroup Robustness Benchmarks. Next, we consider two benchmark subgroup robustness datasets: MultiNLI (multinli) and CelebA (celeba). MultiNLI consists of sentence pairs, where the task is to classify the relationship of the second sentence to the first as entailment, neutral, or contradiction. The spurious feature is the presence of negation words, which often indicate contradiction. CelebA contains images of celebrity faces, with the task of predicting whether the hair color is blonde/not blonde, and the celebrity’s gender (male/female) serves as the spurious attribute.
@@
We fine-tune a pretrained BERT bert-base-uncased model (devlin-bert) on the MultiNLI dataset, and an ImageNet-pretrained ResNet-50 model on CelebA, using different optimizers and evaluate the average (group-balanced) test accuracy and the worst-group accuracy over the number of finetuning epochs. To ensure our comparisons are robust, we conduct extensive hyperparameter sweeps for all optimizers (see Sec.˜D.4 for details). We select the best hyperparameter configuration for any given optimizer and number of fine-tuning epochs, $T$, based on the highest average worst-group accuracy achieved on the validation set, averaged across multiple initialization seeds, at any point up to $T$ epochs. We then report the performance for the optimal configuration for different FT epochs. From Fig.˜8, we see that Muon outperforms SGD on both datasets. On MultiNLI, Adam attains better performance than Muon/Shampoo whereas on CelebA, its performance is comparable to Muon. These experiments show that Muon promotes more balanced learning of spectral components than SGD, thereby improving generalization.
@@
Here, we extend our investigation of class imbalance to language modeling. While our previous experiments and theory focused on one-hot classification, we now shift our comparison of optimizers to a next-token prediction (NTP) task. This task presents a natural imbalance over the distribution of tokens, as word frequencies follow a long-tail distribution described by Zipf’s law (Piantadosi2014ZipfsWF).
@@
TinyStories Dataset. To examine how different optimizers compare on rare token learning, we train small Transformer models on the TinyStories (eldan-li-2023-tinystories) corpus and compare the performance of Muon, SGD, and Adam. We group tokens into "frequent" and "rare" categories based on their corpus statistics and evaluate the token-level accuracy for each group separately on a held-out test set. Complete training details are available in App. D.5. The performance on rare and frequent tokens is reported in Fig.˜9. Consistent with our findings in the one-hot class-imbalanced classification, we observe that Muon generalizes faster on both the majority (frequent) and minority (rare) components compared to SGD, and in this case, even compared to Adam.
Attribute-Organism Classification. To extend our findings from statistical imbalance to a setting with explicit conceptual structure, we conduct a toy experiment on hierarchical concept learning (zhao2025geometrysemanticsnexttokenprediction). In this task, a model must classify organisms (e.g., "dog") based on their attributes, which are shared at different levels of a taxonomic tree (e.g., "needs oxygen" for all animals). We observe distinct learning dynamics: GD learns the coarse, high-level categories (plant vs. animal) first, while Muon exhibits a more balanced learning progression across both coarse and fine-grained concepts (see Sec.˜D.6 for details)
@@ 5 Discussion and Outlook
A crucial test of this insight is how its practical variants, Muon and Shampoo, perform against strong, widely-used baselines like Adam. In our experiments in Sec.˜4, we find that Adam’s generalization performance is only slightly worse—or sometimes even better—than that of Muon and Shampoo. This is consistent with recent large-scale benchmarks for LLM pretraining, which show that well-tuned Adam variants remain highly competitive (wen2025fantasticpretrainingoptimizers; semenov2025benchmarkingoptimizerslargelanguage). Previous work has used different perspectives to explain Adam’s superiority over (S)GD: i) kunstner2024heavy show how heavy-tailed class-imbalanced settings lead to gains for Adam by focusing on their optimization dynamics; ii) vasudeva2025richsimpleimplicitbias show that Adam promotes richer feature learning than SGD in neural networks, which can lead to better generalization in group-imbalanced settings. Together, these works underscore Adam’s advantages in imbalanced regimes, albeit through different mechanisms. Our motivation for using this testbed stems from recognizing that spectral components of imbalanced data directly correspond to majority and minority classes or groups, and our analysis reveals that SpecGD achieves its gains through a fundamentally different and interpretable mechanism—balanced learning of principal components.
A natural question is how the implicit balancing effect of spectral methods compares to explicitly designed techniques like loss re-weighting (byrd2019effect; TengyuMa; sagawa2019distributionally; Menon; pmlr-v139-liu21f). In Sec.˜D.8, we show that while weighted cross-entropy (wCE) effectively encourages balanced learning of spectral components on NMD, Muon provides a similar, though weaker, effect "for free" without requiring explicit group priors. This suggests spectral methods could be useful when such information is unavailable for re-weighting.
Our work provides a framework to understand the generalization of spectrum-aware optimizers within a controlled setting using population statistics and squared loss. A natural next step is to build upon this foundation by extending the analysis to more complex regimes using finite samples and cross-entropy (CE) loss. To bridge this gap, we conduct an initial empirical analysis of rate of learning spectral components for nonlinear models trained with CE loss, by tracking singular values of the logit matrices during training (see Sec.˜D.10 for details). We outline future directions and their associated challenges in App.˜F.
@@ Appendix A Optimizers
[…]
Then, the update is $\mathbf{\Delta}_{t}=(\hat{\bm{Z}}_{t}+\epsilon\bm{1}\bm{1}^{\top})^{\circ-1/2}\odot\hat{\bm{M}}_{t}$, where $(\cdot)^{\circ}$ denotes the Hadamard power, $\epsilon>0$ is the numerical precision parameter, and $\mathbf{1}$ denotes the all-ones vector. It is easy to see that Adam reduces to SignGD when we set $\beta_{1}=\beta_{2}=\epsilon=0$.
@@ B.2 Proof of Proposition˜1
We first state the GF equation
$\frac{d}{dt}\bm{W}(t)=-\nabla\mathcal{L}(\bm{W}(t)).$
Using this, we can write the evolution of $\overline{\bm{W}}(t):={\bm{U}}^{\top}\bm{W}(t){\bm{V}}$ as
@@
We first state SpecGF equation
$\frac{d}{dt}\bm{W}(t)=-{\bm{U}}(t){\bm{V}}(t)^{\top},$
where the gradient has SVD $\nabla\mathcal{L}(\bm{W}(t))={\bm{U}}(t)\bm{\Sigma}(t){\bm{V}}(t)^{\top}$. Following the same steps as for SpecGD, we can show that ${\bm{U}}(t)={\bm{U}}$, ${\bm{V}}(t)={\bm{V}}$, and $\overline{\bm{W}}(t)$ evolves as